# VIST3A: Text-to-3D by Stitching a Multi-view Reconstruction Network to a Video Generator

### v1

@@ C.2 Detailed Evaluation Protocol
Details for 3D generation evaluation.
For T3Bench, we evaluate on all 300 prompts, in contrast to prior works that considered only the 100 single-object-with-surroundings subset. SceneBench is evaluated on 80 prompts from the Prometheus3D (Yang et al., 2025c) prompt set, targeting scene-level generation. For DPG-Bench, we sample 100 prompts from the original 1K-prompt dataset.
For Matrix3D-omni, we used their official code for text-to-generation and employed Panorama LRM for reconstruction during inference. For SDS-based methods like SplatFlow and Director3D that perform refinement, we evaluated the final results after SDS optimization. We generate 13 frames for all models using 80 denoising steps, and apply classifier-free guidance (Ho & Salimans, 2022) with a scale of 7.5. We observed that the Gaussian splatting produced by the MVDUSt3R model does not generalize well across diverse domains, often failing to estimate the scale of primitives. To address this issue, we refined the Gaussian primitives using the source view for 100 optimization steps, minimizing a reconstruction loss defined as MSE + 0.05 × LPIPS. For this refinement, we used the Adam optimizer with separate learning rates for each parameter group: 2e-4 for means, 5e-4 for opacity, 5e-4 for scale, 1e-4 for rotation, and 0 for rgbs. This lightweight refinement effectively corrected the scale estimation errors. For our text-to-3DGS evaluation, we render 8 random viewpoints from the generated Gaussian Splatting representations for assessment.
@@
For a more comprehensive assessment of generative quality, we utilize the Unified Reward model (Wang et al., 2025c), which is based on the powerful Qwen 2.5-7B Vision Language Model (Team, 2025) [^2] https://huggingface.co/CodeGoat24/UnifiedReward-qwen-7b. This provides fine-grained, pointwise scores on complex attributes equipped with a powerful understanding capability. By feeding the input prompt and rendered images into a format adapted from the official implementation script [^3] https://github.com/CodeGoat24/UnifiedReward/blob/main/inference_qwen/image_generation/qwen_point_score_ACS_image_generation.py, we obtain scores for three key aspects:
@@
Details for model stitching evaluation.
For novel-view synthesis, we follow prior works (Go et al., 2025a; b) and adopt an 8-frame input setup to evaluate performance on 4 target views. To accommodate the fixed-length input requirements of video VAE architectures due to temporal compression, we pad shorter sequences by duplicating the final frame. For estimating the camera poses of the target views, we adopt the strategy from AnySplat (Jiang et al., 2025), which jointly predicts the poses and renders the corresponding images. This contrasts with previous VAE-based methods that presume access to ground-truth camera poses for rendering.
For pointmap and camera pose estimation evaluation, we use a 13-frame input setup. Since our stitched VAE’s encoder is a video VAE, we arrange the multi-view images (typically provided unordered by previous works) into sequences with smooth view transitions to resemble video input. We adopt Pi3 (Wang et al., 2025d) official evaluation code and follow their preprocessing pipeline.
@@ D.1 Impact of Direct Reward Finetuning
Appendix E Additional Experimental Results
@@
Analysis on searched stitching index.
In Section 4.4, we showed that earlier layers in the network tend to be more linearly correspondent. We extend this analysis to various VAE architectures, including CogVideoX, SVD, Hunyuan, and Wan, paired with MVDUSt3R and AnySplat, to observe the generalizability of this finding.
Additional qualitative results.
We present additional qualitative results of VIST3A with Wan + AnySplat in Fig. 810. Text-to-pointmap generation results obtained by combining VGGT with Wan through VIST3A are shown in Fig. 11. Finally, Fig. 12 illustrates VIST3A results with MVDust3R + Wan.

### v2

@@ C.2 Detailed Evaluation Protocol
Details for 3D generation evaluation. For T3Bench, we evaluate on all 300 prompts, in contrast to prior works that considered only the 100 single-object-with-surroundings subset. SceneBench is evaluated on 80 prompts from the Prometheus3D (Yang et al., 2025c) prompt set, targeting scene-level generation. For DPG-Bench, we sample 100 prompts from the original 1K-prompt dataset.
For Matrix3D-omni, we used their official code for text-to-generation and employed Panorama LRM for reconstruction during inference. For SDS-based methods like SplatFlow and Director3D that perform refinement, we evaluated the final results after SDS optimization. We generate 13 frames for all models using 80 denoising steps, and apply classifier-free guidance (Ho and Salimans, 2021) with a scale of 7.5. We observed that the Gaussian splatting produced by the MVDUSt3R model does not generalize well across diverse domains, often failing to estimate the scale of primitives. To address this issue, we refined the Gaussian primitives using the source view for 100 optimization steps, minimizing a reconstruction loss defined as MSE + 0.05 × LPIPS. For this refinement, we used the Adam optimizer with separate learning rates for each parameter group: 2e-4 for means, 5e-4 for opacity, 5e-4 for scale, 1e-4 for rotation, and 0 for rgbs. This lightweight refinement effectively corrected the scale estimation errors. For our text-to-3DGS evaluation, we render 8 random viewpoints from the generated Gaussian Splatting representations for assessment.
@@
For a more comprehensive assessment of generative quality, we utilize the Unified Reward model (Wang et al., 2025d), which is based on the powerful Qwen 2.5-7B Vision Language Model (Team, 2025) [^2] https://huggingface.co/CodeGoat24/UnifiedReward-qwen-7b. This provides fine-grained, pointwise scores on complex attributes equipped with a powerful understanding capability. By feeding the input prompt and rendered images into a format adapted from the official implementation script [^3] https://github.com/CodeGoat24/UnifiedReward/blob/main/inference_qwen/image_generation/qwen_point_score_ACS_image_generation.py, we obtain scores for three key aspects:
@@
Details for model stitching evaluation. For novel-view synthesis, we follow prior works (Go et al., 2025a; b) and adopt an 8-frame input setup to evaluate performance on 4 target views. To accommodate the fixed-length input requirements of video VAE architectures due to temporal compression, we pad shorter sequences by duplicating the final frame. For estimating the camera poses of the target views, we adopt the strategy from AnySplat (Jiang et al., 2025), which jointly predicts the poses and renders the corresponding images. This contrasts with previous VAE-based methods that presume access to ground-truth camera poses for rendering.
For pointmap and camera pose estimation evaluation, we use a 13-frame input setup. Since our stitched VAE’s encoder is a video VAE, we arrange the multi-view images (typically provided unordered by previous works) into sequences with smooth view transitions to resemble video input. We adopt Pi3 (Wang et al., 2026) official evaluation code and follow their preprocessing pipeline.
C.3 Selection Criterion for 3D models
In this section, we elaborate on the criteria used for selecting 3D foundation models in VIST3A.
Our primary criterion for selecting a 3D backbone is the scale of the pretraining dataset, which mainly determines the generalizability of the model. This is because the coupled 3D model should cover highly diverse domains of video generative models. AnySplat (Jiang et al., 2025) is selected as it represents one of the few Gaussian Splatting models trained on such a scale. For pointmap-based models, we utilize MVDUSt3R (Wang et al., 2024) for its balance of dataset scale and efficiency (facilitating feasibility checks), and VGGT (Wang et al., 2025b) for its superior performance and pretraining scale.
@@ D.1 Impact of Direct Reward Finetuning
Figure 7: Qualitative comparison of different finetuning strategies. Pretrained video model (No-Finetuning) produces dynamic videos, causing severe ghosting in the reconstructed 3D scenes. Multi-view finetuning (Multi-view only) reduces this motion and improves 3D consistency but introduces semantic and quality degradations, while our direct reward tuning (Ours) yields sharper renderings that better align with the input text prompts.
@@
To further analyze the effect of our training strategy, we provide a qualitative comparison between the pretrained video model, the multi-view-only finetuning baseline, and our reward-tuned model in Fig. 7. As shown in the decoded frames, the No-Finetuning baseline (pretrained video generator) often produces dynamic videos with temporal motion. This leads to severe ghosting artifacts in the generated 3D scenes, clearly visible as multiple outlines of the footballs and the car.
The Multi-view only baseline (second row) effectively suppresses this motion, enforcing 3D consistency. However, relying solely on a multi-view dataset limits the model’s generalizability, and the lack of alignment with the decoder hinders the model from generating latents that are well reconstructible by the decoder. As a result, the generations exhibit semantic and quality degradations: the “green” football is missing in the first example, and the car becomes blurrier in the second example.
In contrast, our direct reward tuning (Ours) applies rewards to the decoded 3D representation of the stitched decoder, explicitly encouraging high reconstruction quality, 3D consistency, and better alignment with the text prompt. Consequently, it produces sharper and more faithful renderings than the baselines in both examples.
@@
Appendix E Additional Results
@@
Figure 9: Viewpoint control in 3DGS generation with prompts. (Top) The prompt containing “Aerial dronshot” results in a high-angle downward perspective. (Bottom) The prompt containing “Camera pans left to right” generates a trajectory where the viewpoint sweeps across the scene from left to right.
Prompt-based camera control. Modern video generative model, including Wan 2.1, can reflect camera-related prompts such as “aerial view” and “camera pans left to right” in the generated videos. As our framework is built upon the Wan 2.1 backbone, we inherit this property, enabling text-driven viewpoint control in 3D generation. As illustrated in Fig. 9, the prompt containing “Aerial dronshot” induces a high-angle, downward-looking perspective (Top), whereas the prompt with “Camera pans left to right” results in a horizontal sweeping motion that traverses the scene from left to right (Bottom). This shows that our model effectively inherits the semantic camera instruction understanding capability of pretrained video generative models.
Table 7: Video generation performance comparison between the original video generator and VIST3A on VBench.
Model Quality Score Semantic Score Total Score
Original video generator 0.7827 0.7275 0.7716
VIST3A (Wan + AnySplat) 0.8143 0.7143 0.7943
Impact of VIST3A finetuning on video generation performance. We aim for video generative models to produce 3D-consistent frames, and we propose a finetuning strategy toward this goal. Here, we evaluate how much this finetuning degrades the original model’s generative capability. To measure this, we use VBench (Huang et al., 2024), a well-established benchmark for video generation. We generate videos using the original model by matching the resolution to the VIST3A inference setting, and compare the video results of the finetuned model with VIST3A with Wan and AnySplat.
As shown in Table 7, the VIST3A-adapted model achieves a higher quality score and a higher total score than the original video generator, with only a marginal difference in the Semantic Score. This demonstrates that VIST3A’s finetuning does not significantly degrade the video model’s generative capability—in fact, the overall VBench performance is slightly improved under the VIST3A setup.
Analysis on searched stitching index. In Section 4.4, we showed that earlier layers in the network tend to be more linearly correspondent. We extend this analysis to various VAE architectures, including CogVideoX, SVD, Hunyuan, and Wan, paired with MVDUSt3R and AnySplat, to observe the generalizability of this finding.
@@
Additional qualitative results. We present additional qualitative results of VIST3A with Wan + AnySplat in Fig. 1113. Text-to-pointmap generation results obtained by combining VGGT with Wan through VIST3A are shown in Fig. 14. Finally, Fig. 15 illustrates VIST3A results with MVDust3R + Wan.